OpenStudy (rayman5766):

The triangles shown are similar. What is the value of x? VHS_PA_S2_01_L210_L310_Q8.gif A. 17 B. 5 C. 10 D. 20

OpenStudy (rayman5766):

@texaschic101

OpenStudy (texaschic101):

I cannot see the picture

OpenStudy (rayman5766):

cant you click on it I thought you could hem hold on

OpenStudy (rayman5766):

yep how did made blue like that

OpenStudy (texaschic101):

copy and paste.

OpenStudy (rayman5766):

I did

OpenStudy (rayman5766):

I think its D

OpenStudy (texaschic101):

you can set this up as a proportion 4/16 = 5/x --- cross multiply 4x = 80 x = 20 you think correctly...lol

OpenStudy (rayman5766):

yea